Methane, the main component of natural gas, burns as follows. Accordingly, how many kJ heat is released from the full combustion of 1.63x103L methane at 18.60C and 768mmHg pressure? Also, considering that this reaction occurs in 100% efficiency, how many liters of water can increase the temperature from 8.70C to 60.00C if used in heating?
1
Expert's answer
2020-05-06T13:58:12-0400

1.CH4 + O2 --> CO2+H2O;

∆H=802.3 kJ/mol


p=768 mm Hg=102,4 kPa

T=18.6 oC=273+18.6=291.6 K

n(CH4)=(p*V)/(R*T)=102.4*1.63*(10^3)/(8.314*296.7)=67.66 mol


Q=∆H*n=802.3*67.66=54283.62 kJ


2.C(H2O)=4.2 kJ/(kg*oC)

C=Q/(m*∆T);

m(H2O)=Q/(C*∆T)=54283.62/(4.2*51.3)=252 kg

L(H2O)=252 L

Answer: 1. 54283.62 kJ

2.252 L